Waiting period. Upstate University currently has a 6 comma 200​-car parking capacity for​ faculty, staff, and students. This​ year, the university issued 4,424 parking passes. The number of parking passes issued has been growing at a rate of 6​% per year. How long will it be before the university will need to add parking​ spaces?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

6200=4424*(1.06)^n

(6200/4424)=(1.06)^n

Taking log on both sides;

log(6200/4424)=n*log(1.06)

n=log(6200/4424)/log(1.06)

=5.79 years(Approx).